OpenStudy (bimbim2697):

solving system using subsitution question a shopper purchased 8 t-shirts and 5 pairs of pants for $220.The next day he purh pair of purchased 5 tshirts and 1 pair of pants for $112.How much does each tshirt and each pair of pants cost?

OpenStudy (anonymous):

8t + 5p = 220 5t + p = 112 From the second equation: p = 112 - 5t Replace into first: 8t + 5(112 - 5t) = 220 t = 20 p = 112 - 5t = 112 - 100 = 12

OpenStudy (bimbim2697):

thank u so much